OpenStudy (anonymous):

Let f(x)=5/x and g (x)=2x^2+5x ,What two numbers are not in the domain of f o g?

OpenStudy (nurali):

The two zeros of g(x) will feed 0 into f(g(x)) = f o g(x), and f(0) = 5/0 is undefined. So the answer is the two zeros of g(x) = 2x^2 + 5x = 0 which are x = 0, -5/2

OpenStudy (luigi0210):

since g(x) is the domain of the new function we have to set it equal to zero and solve

OpenStudy (luigi0210):

wait.. which is NOT?